# Aspects of the commuting graph [^1] To the memory of Richard Parker

### v3

@@ 10 Preliminaries
We conclude this section with a general remark. In the commuting graph $\Gamma(G)$ of a group $G$, elements of the center $Z(G)$ are dominating vertices; this is the reason why, to investigate questions about connectedness, they are excluded. However, in our case, it doesn’t matter:
@@
11 A structure of groups whose commuting graphs are cographs or chordal graphs
@@
First, we investigate special cases of groups like direct and central products and Frobenuis groups whose commuting graph is a cograph or a chordal graph, and then prove the general structure theorem.
@@ 11.1 Direct and central products
By Lemma 11.1 and Observation 3, the following result holds.
@@
is a $C_{4}$ subgraph which projects isomorphically to $H\circ K$. $\qed$
@@ 11.2 Frobenius groups
In this sections we investigate when the commuting graph of a Frobenius group is a cograph or a chordal graph. We prove the following assertion.
@@
Since $K$ is nilpotent, Corollary 11.3 gives a criterion when $\Gamma(K)$ is a cograph (resp. a chordal graph).
@@ 11.3 A general structure theorem
Let $G$ be a finite group such that $\Gamma(G)$ is a cograph (resp. a chordal graph), and let $F(G)$ be the Fitting subgroup of $G$ (the largest nilpotent normal subgroup of $G$).
@@
Proof 20
Let $F^{*}(G)$ be the generalized Fitting subgroup of $G$. By [3, 31.12],
$F^{*}(G)=F(G)\circ E(G),$ (8)
where $E(G)$ is the subgroup of $G$ generated by the set $\mathop{\mathrm{Comp}}(G)$ of all components (subnormal quasisimple subgroups) of $G$.
Let $H_{1}$ and $H_{2}$ be distinct elements from $\mathop{\mathrm{Comp}}(G)$. By [3, 31.5], $H_{1}$ and $H_{2}$ commute, therefore $\langle H_{1},H_{2}\rangle=H_{1}\circ H_{2}$ is a subgroup of $G$. By Lemma 11.4, $\Gamma(G)$ is neither a cograph nor a chordal graph, a contradiction to Proposition 10.2. Thus, $|\mathop{\mathrm{Comp}}(G)|\leq 1$.
@@
Now we prove statement (b). Assume that $L=\mathop{\mathrm{Soc}}(G)$ is non-abelian. Then $L=A\times S$, where $A$ is abelian and $S$ is the direct product of non-abelian simple groups, moreover, both $A$ and $S$ are normal subgroups of $G$. By Propositions 10.2 and 11.2, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph). If $C=C_{G}(S)$, then $C\unlhd G$ and $C\cap S=1$ since $S$ is simple. Thus, $N=SC=S\times C$ is a normal subgroup of $G$. By Proposition 10.2, $\Gamma(N)$ is also a cograph (resp. a chordal graph). Using Proposition 11.2, we conclude that $C$ is abelian. Moreover, $G/C$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(S)$ which is almost simple, therefore $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$. Thus, by the Schreier Conjecture, $G/N$ is solvable, and therefore statement (b) holds. $\qed$
@@ 12.2 Generalized dihedral groups and generalized quaternion groups
Let $Q_{4m}$ be the generalized quaternion group as presented in Equationi (4). Then $Q_{4m}$ is an AC-group. Hence $\Gamma(Q_{4m})$ is a cograph as well as a chordal graph.
Proof 24
The proof is similar to the preceding theorem. We find that, with $G=Q_{4m}$, we have $Z(G)=\{e,x^{m}\}$; for $11\leq i\leq 2m-$ 1 and $i\neq m$, we find that $C_{G}(x^{i})=\langle x\rangle$, while for $10\leq i\leq 2m-$ 1 we have $C_{G}(x^{i}y)=\langle x^{i}y\rangle=\{e,x^{i}y,x^{m},x^{m+i}y\}$. So again the induced subgraph on $G\setminus Z(G)$ consists of a complete graph of size $22m-$ 2 together with $m$ disjoint edges. $\qed$
@@ 12.3 Symmetric groups and alternating groups
The symmetric group, denoted by $\mathop{\mathrm{Sym}}\nolimits(n)$ or $S_{n}$, is the group of permutations on $n$ letters. As we known, the symmetric group is important in many different areas of mathematics, including combinatorics and group theory. Note that $S_{n}$ is abelian if and only if $n\leq 2$.
@@
Proof 25
Proposition cog-subg shows that it is enough to show that $\Gamma(S_{3})$ is a cograph but $\Gamma(S_{4})$ is not (since $S_{n-1}\leq S_{n}$ for any $n$). By Proposition 12.5, we have that $\Gamma(S_{3})$ is a cograph as $S_{3}\cong D_{6}$. i As $S_{n-1}\leq S_{n}$, Proposition 10.2 shows that if $n\leq 3$, then $\Gamma(S_{n})$ is a cograph. In $\Gamma(S_{4})$,
@@
Suppose, for the sake of contradiction, that $\Gamma(S_{4})$ contains an induced subgraph isomorphic to $C_{n}$ with $n\geq 4$, say $\Delta$. Let $a,b,c,d$ be $4$ distinct vertices of $\Delta$. Then $C_{S_{4}}(x)$ is non-abelian for any $x\in\{a,b,c,d\}$. Note that every centralizer in 10 is abelian. It follows from 9 that every element in $\{a,b,c,d\}$ has cycle type $(ij)(kl)$, which is impossible because $S_{4}$ has precisely three elements that are the product of two disjoint cycles. $\qed$
The group consisting of all even permutations on a set of length $n$ is called the alternating group and is denoted by $\mathop{\mathrm{Alt}}\nolimits(n)$ or $A_{n}$. Note that $A_{n}$ is a simple group for each $n\geq 5$.
@@
Proof 27
The group $A_{5}$ is a CA-group (it is easily checked that all its non-abelian subgroups have trivial centre), so the result for $n=5$ follows from Proposition 12.2. Moreover, $A_{n}\leq A_{5}$ for $n\leq 5$, so the result follows for these groups as well.
@@ 12.5 Simple groups of Lie type whose commuting graph is a cograph
Suppose next that $q$ is odd at least $5$. Let $\epsilon\in\{+1,-1\}$ be chosen so that $q+\epsilon 1$ is divisible by $4$. Note that $\mathrm{PSL}(2,q)$ contains a Klein group such that the centralizers of its involutions are Dihedral groups $D_{2(q+\epsilon 1)}$ which contain disjoint cyclic groups of order $(q+\epsilon)/2$. If $q>5$, then $(q+\varepsilon)/2>2$. So two involutions together with generators of cyclic groups of order $(q+\epsilon)/2$ in their centralizers induce a path of length $4$. (The list of subgroups shows that these generators do not commute.) So $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is not a cograph. $\qed$
@@
(see the $\mathbb{ATLAS}$ of Finite Groups [28]). So (a), (b) andn (d) are covered by our earlier results, and (c) and (e) will be treated as symplectic or unitary groups.
@@
It suffices to show that the graph obtained by removing the identity is a cograph. Now $\mathop{\mathrm{Sz}}\nolimits(q)$ has maximal cyclic subgroups of orders $q-1$ and $q\pm\sqrt{2q}+1$, and every element of odd order lies in one of these; so the non-identity elements of odd order carry a disjoint union of complete graphs, with no edges to the remaining elements. (The information used in this paragraph and the next is taken from Suzuki’s original paper [72]; it can also be found in [77, Section 4.2].) The remaining elements have order $2$ or $4$, and since $\mathop{\mathrm{Sz}}\nolimits(q)$ is a CIT-group (one in which the centralizer of any involution is a $2$-group), the remaining vertices lie in a disjoint union of copies of the commuting graph on $P\setminus\{e\}$, where $P$ is a Sylow $2$-subgroup. Now $|P|=q^{2}$ and $|Z(P)|=q$; elements of $Z(P)$ are joined to every element of $P$ so we may remove them. The remaining elements of $P$ have order $4$. If $g$ is one of these, then $|C_{P}(g)|=2q$ and $C_{P}(g)$ consists of two cosets of $Z(P)$, so it is abelian; thus the elements of order $4$ fall into $q-1$ sets of size $q$, each set a complete graph with no edges between different sets. Thus the graph is indeed a cograph. $\qed$
@@
The Tits group ${}^{2}F_{4}(2)^{\prime}$ contains a subgroup isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,25)$ (this can be found in the $\mathbb{ATLAS}$ of Finite Groups [28]); so its commuting graph is not a cograph. This group is a subgroup of the large Ree group ${}^{2}F_{4}(2^{d})$ for all odd $d$ [52].
@@ 13 Appendix
Then $Z(G)a$ has order $p$ or $q$. Thus, $C_{G}(a)/Z(G)$ has an element, say $Z(G)b$, such that $C_{G}(a)/Z(G)=\langle Z(G)b,Z(G)a\rangle$. Note that $[b,a]=e$. Hence, $C_{G}(a)=\langle b,a,Z(G)\rangle$ is an abelian group, as desired. $\qed$
@@
    $|G|=pqr$ or $pq$, where $p,q,r$ are primes;
@@
Proof 34
Proposition 12.7 implies that $\Gamma(S_{4})$ is not a cograph. Thus, it suffices to show that if $|G|\leq 24$, then $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$. Now suppose that $|G|\leq 24$. It follows from Lemma 13.3 that if $|G|\notin\{16,24\}$, then $\Gamma(G)$ is a cograph. If $|G|=16$, since a $2$-group has non-trivial center, we have that $|G/Z(G)|=2^{k}$ with $k\leq 3$, and so Lemma 13.3(c) implies that $\Gamma(G)$ is a cograph. Finally, we suppose that $|G|=24$. Note that the fact that if $G\ncong S_{4}$, then $G$ has non-trivial center (cf. Burnside [18]). So, by Lemma 13.3(c), we have that if $G\ncong S_{4}$, then $\Gamma(G)$ is a cograph, as desired. $\qed$
@@
$D_{8}\circ D_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{2}=w^{2}=[x,y]^{2}=[z,w]^{2}=e,$ (12)
$[x,z]=[z,y]=[y,w]=[w,x]=e\rangle$
@@
$D_{8}\circ Q_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{4}=w^{4}=[x,y]^{2}=[z,w]^{2}=e,$ (14)
$[x,z]=[z,y]=[y,w]=[w,x]=e\rangle$
@@
Now, it suffices to show that if $|G|<32$, then $\Gamma(G)$ is a chordal graph. Suppose that $|G|\leq 31$. Note that a $p$-group has non-trivial center where $p$ is a prime. From Lemma 13.3, it follows that if $|G|\neq 24$, then $\Gamma(G)$ is a chordal graph. Now consider $|G|=24$. If $G\cong S_{4}$, then by Proposition 12.8, $\Gamma(G)$ is chordal. If $G\ncong S_{4}$, then by Lemma 13.3 and the proof of Proposition 13.4, $\Gamma(G)$ is chordal, as desired. $\qed$

### v4

@@ 10 Preliminaries
We conclude this section with a general remark. In the commuting graph $\Gamma(G)$ of a group $G$, elements of the centre $Z(G)$ are dominating vertices; this is the reason why, to investigate questions about connectedness, they are excluded. However, in our case, it doesn’t matter:
@@
11 Structure of groups whose commuting graphs are cographs or chordal graphs
@@
First, we investigate special cases of groups like direct and central products and Frobenius groups whose commuting graph is a cograph or a chordal graph, and then prove the general structure theorem.
@@ 11.1 Direct and central products
By Lemma 11.1 and Observation 4, the following result holds.
@@
is a subgraph isomorphic to $C_{4}$ which projects isomorphically to $H\circ K$. $\qed$
@@ 11.2 Frobenius groups
In this sections we investigate when the commuting graph of a Frobenius group is a cograph or a chordal graph. We use the structure of the commuting graphs of Frobenius groups established in Theorem 6.2. We prove the following assertion.
@@
Since $K$ is nilpotent, Corollary 11.3 gives a criterion for $\Gamma(K)$ to be a cograph (resp. a chordal graph).
@@ 11.3 A general structure theorem
Let $G$ be a finite group such that $\Gamma(G)$ is a cograph (resp. a chordal graph), and let $F(G)$ be the Fitting subgroup of $G$.
@@
Proof 28
Let $F^{*}(G)$ be the generalized Fitting subgroup of $G$. Let $H_{1}$ and $H_{2}$ be distinct elements from $\mathop{\mathrm{Comp}}(G)$. By [3, 31.5], $H_{1}$ and $H_{2}$ commute, therefore $\langle H_{1},H_{2}\rangle=H_{1}\circ H_{2}$ is a subgroup of $G$. By Lemma 11.4, $\Gamma(G)$ is neither a cograph nor a chordal graph, a contradiction to Proposition 10.2. Thus, $|\mathop{\mathrm{Comp}}(G)|\leq 1$.
@@
Now we prove statement (b). Assume that $L=\mathop{\mathrm{Soc}}(G)$ is non-abelian. Then $L=A\times S$, where $A$ is abelian and $S$ is the direct product of non-abelian simple groups, moreover, both $A$ and $S$ are normal subgroups of $G$. By Propositions 10.2 and 11.2, $S$ is simple and $\Gamma(S)$ is a cograph (resp. a chordal graph). If $C=C_{G}(S)$, then $C\unlhd G$ and $C\cap S=1$ since $S$ is simple. Thus, $N=SC=S\times C$ is a normal subgroup of $G$. By Proposition 10.2, $\Gamma(N)$ is also a cograph (resp. a chordal graph). Using Proposition 11.2, we conclude that $C$ is abelian. Moreover, $G/C$ is isomorphic to a subgroup of $\mathop{\mathrm{Aut}}(S)$ which is almost simple, therefore $G/N$ is isomorphic to a subgroup of $\mathop{\mathrm{Out}}(S)$. Thus, by the Schreier Conjecture (Theorem 2.1), $G/N$ is solvable, and therefore statement (b) holds. $\qed$
@@ 12.2 Generalized dihedral groups and generalized quaternion groups
Let $Q_{4m}$ be the generalized quaternion group as presented in Equation (5). Then $Q_{4m}$ is an AC-group. Hence $\Gamma(Q_{4m})$ is a cograph as well as a chordal graph.
Proof 32
The proof is similar to the preceding theorem. We find that, with $G=Q_{4m}$, we have $Z(G)=\{e,x^{m}\}$ (Observation 2); for $11\leq i\leq 2m-$ 1 and $i\neq m$, we find that $C_{G}(x^{i})=\langle x\rangle$, while for $10\leq i\leq 2m-$ 1 we have $C_{G}(x^{i}y)=\langle x^{i}y\rangle=\{e,x^{i}y,x^{m},x^{m+i}y\}$. So again the induced subgraph on $G\setminus Z(G)$ consists of a complete graph of size $22m-$ 2 together with $m$ disjoint edges. $\qed$
@@ 12.3 Symmetric groups and alternating groups
Proof 33
Proposition 10.2 shows that it is enough to show that $\Gamma(S_{3})$ is a cograph but $\Gamma(S_{4})$ is not (since $S_{n-1}\leq S_{n}$ for any $n$). By Proposition 12.5, we have that $\Gamma(S_{3})$ is a cograph as $S_{3}\cong D_{6}$. As $S_{n-1}\leq S_{n}$, Proposition 10.2 shows that if $n\leq 3$, then $\Gamma(S_{n})$ is a cograph. In $\Gamma(S_{4})$,
@@
Suppose, for the sake of contradiction, that $\Gamma(S_{4})$ contains an induced subgraph isomorphic to $C_{n}$ with $n\geq 4$, say $\Delta$. Let $a,b,c,d$ be $4$ distinct vertices of $\Delta$. Then $C_{S_{4}}(x)$ is non-abelian for any $x\in\{a,b,c,d\}$. Note that every centralizer in (11) is abelian. It follows from (10) that every element in $\{a,b,c,d\}$ has cycle type $(ij)(kl)$, which is impossible because $S_{4}$ has precisely three elements that are the product of two disjoint cycles. $\qed$
@@
Proof 35
The group $A_{5}$ is a CA-group (it is easily checked from [28] that all its non-abelian subgroups have trivial centre), so the result for $n=5$ follows from Proposition 12.2. Moreover, $A_{n}\leq A_{5}$ for $n\leq 5$, so the result follows for these groups as well.
@@ 12.5 Simple groups of Lie type whose commuting graph is a cograph
Suppose next that $q$ is odd at least $5$. Let $\epsilon\in\{+1,-1\}$ be chosen so that $q+\epsilon 1$ is divisible by $4$. Note that $\mathrm{PSL}(2,q)$ contains a Klein group such that the centralizers of its involutions are Dihedral groups $D_{2(q+\epsilon 1)}$ which contain disjoint cyclic groups of order $(q+\epsilon)/2$. If $q>5$, then $(q+\varepsilon)/2>2$. So two involutions together with generators of cyclic groups of order $(q+\epsilon)/2$ in their centralizers induce a path of length $4$. (The list of subgroups of $G$ shows that these generators do not commute.) So $\Gamma(\mathop{\mathrm{PSL}}\nolimits(2,q))$ is not a cograph. $\qed$
@@
(see the $\mathbb{ATLAS}$ of Finite Groups [28]). So (a), (b) and (d) are covered by our earlier results, and (c) and (e) will be treated as symplectic or unitary groups.
@@
The information used in the next two paragraphs is taken from Suzuki’s original paper [74]; it can also be found in [79, Section 4.2].
It suffices to show that the graph obtained by removing the identity is a cograph. Now $\mathop{\mathrm{Sz}}\nolimits(q)$ has maximal cyclic subgroups of orders $q-1$ and $q\pm\sqrt{2q}+1$, and every element of odd order lies in one of these; so the non-identity elements of odd order carry a disjoint union of complete graphs, with no edges to the remaining elements. The remaining elements have order $2$ or $4$, and since $\mathop{\mathrm{Sz}}\nolimits(q)$ is a CIT-group (one in which the centralizer of any involution is a $2$-group), the remaining vertices lie in a disjoint union of copies of the commuting graph on $P\setminus\{e\}$, where $P$ is a Sylow $2$-subgroup. Now $|P|=q^{2}$ and $|Z(P)|=q$; elements of $Z(P)$ are joined to every element of $P$ so we may remove them. The remaining elements of $P$ have order $4$. If $g$ is one of these, then $|C_{P}(g)|=2q$ and $C_{P}(g)$ consists of two cosets of $Z(P)$, so it is abelian; thus the elements of order $4$ fall into $q-1$ sets of size $q$, each set a complete graph with no edges between different sets. Thus the graph is indeed a cograph. $\qed$
@@
The Tits group ${}^{2}F_{4}(2)^{\prime}$ contains a subgroup isomorphic to $\mathop{\mathrm{PSL}}\nolimits(2,25)$ (this can be found in the $\mathbb{ATLAS}$ [28]); so its commuting graph is not a cograph. The group ${}^{2}F_{4}(2)^{\prime}$ is a subgroup of the large Ree group ${}^{2}F_{4}(2^{d})$ for all odd $d$ [54].
@@ 13 Appendix
Then $Z(G)a$ has order $p$ or $q$. Thus, $C_{G}(a)/Z(G)$ has an element, say $Z(G)b$, such that $C_{G}(a)/Z(G)=\langle Z(G)b,Z(G)a\rangle$. Note that $[b,a]=1$. Hence, $C_{G}(a)=\langle b,a,Z(G)\rangle$ is an abelian group, as desired. $\qed$
@@
    $|G|=pqr$ or $pq$, where $p,q,r$ are primes (not necessarily distinct);
@@
Proof 42
Proposition 12.7 implies that $\Gamma(S_{4})$ is not a cograph. Thus, it suffices to show that if $|G|\leq 24$, then $\Gamma(G)$ is a cograph if and only if $G\ncong S_{4}$. Now suppose that $|G|\leq 24$. It follows from Lemma 13.3 that if $|G|\notin\{16,24\}$, then $\Gamma(G)$ is a cograph. If $|G|=16$, since a $2$-group has non-trivial centre, we have that $|G/Z(G)|=2^{k}$ with $k\leq 3$, and so Lemma 13.3(c) implies that $\Gamma(G)$ is a cograph. Finally, we suppose that $|G|=24$. Note that the fact that if $G\ncong S_{4}$, then $G$ has non-trivial centre (cf. Burnside [18]). So, by Lemma 13.3(c), we have that if $G\ncong S_{4}$, then $\Gamma(G)$ is a cograph, as desired. $\qed$
@@
$D_{8}\circ D_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{2}=w^{2}=[x,y]^{2}=[z,w]^{2}=1,$ (13)
$[x,z]=[z,y]=[y,w]=[w,x]=1\rangle$
@@
$D_{8}\circ Q_{8}$ $=$ $\langle x,y,z,w|x^{2}=y^{2}=z^{4}=w^{4}=[x,y]^{2}=[z,w]^{2}=1,$ (15)
$[x,z]=[z,y]=[y,w]=[w,x]=1\rangle$
@@
Now, it suffices to show that if $|G|<32$, then $\Gamma(G)$ is a chordal graph. Suppose that $|G|\leq 31$. Note that a $p$-group has non-trivial centre where $p$ is a prime. From Lemma 13.3, it follows that if $|G|\neq 24$, then $\Gamma(G)$ is a chordal graph. Now consider $|G|=24$. If $G\cong S_{4}$, then by Proposition 12.8, $\Gamma(G)$ is chordal. If $G\ncong S_{4}$, then by Lemma 13.3 and the proof of Proposition 13.4, $\Gamma(G)$ is chordal, as desired. $\qed$
@@ Acknowledgements
Acknowledgement.  We are grateful to the referee for thoughtful comments and suggestions that have helped improve our presentation.